OpenStudy (anonymous):

Can someone help me find the slope and y intercept please? I'd really appreciate it. y= -1/3x + 3 -x + 3y = 9

OpenStudy (mathstudent55):

The slope-intercept form of the equation of a line is y = mx + b where m = slope, and b = y-intercept

OpenStudy (shaik0124):

slope is-1/3 n y intercept is 3

OpenStudy (ash2326):

Standard form of a line with slope m and y-intercept c \[y=mx+c\] Let's compare this with the first equation \[y=-\frac 1 3 x+3\] Do you see any correspondence between the two equations? @unheard

OpenStudy (anonymous):

Kind of, @ash2326 I'm honestly terrible at slope.

OpenStudy (ash2326):

@shaik0124 Please don't give just answers, it won't help the user. Please explain how you got it

OpenStudy (shaik0124):

just compare two equations where m is slope and c is y intrcpt

OpenStudy (shaik0124):

ok aishwarya2326

OpenStudy (anonymous):

I think I understand the comparing the equations.

OpenStudy (shaik0124):

good now corelate both n find values

OpenStudy (anonymous):

I don't understand that o.o

OpenStudy (ash2326):

\[y=mx+c\] \[y=-\frac 13 x+3\] if in the first equation, we put m=-1/3 and c=3 We'll get the second one, won't we?

OpenStudy (anonymous):

Yes. @ash2326

OpenStudy (ash2326):

So there you have m= slope =-1/3 and y intercept c= 3 That's comparing the two equations

OpenStudy (anonymous):

Awesome, @ash2326 . Thanks so much. Could you help me with two more?

OpenStudy (ash2326):

Yeah, I could. If they are similar, you should try them first

OpenStudy (anonymous):

No, the whole thing is part of a small project, but I am no good in Math, excpecially in slopes, haha. How would I graph the line in slope- intercept method? I just need to describe it.

OpenStudy (ash2326):

OK, I'll help you,but first close this question and post it as a new question. That way more users can help

OpenStudy (anonymous):

Awesome, okay.